# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v3

@@ Abstract
Money laundering presents a pervasive challenge, burdening society by financing illegal activities. The use of network information is increasingly being explored to more effectively combat money laundering, given it involves connected parties. This led to a surge in research on network analytics (NA) for anti-money laundering (AML). The literature on NA for AML is, however, fragmented and a comprehensive overview of existing work is missing. This results in limited understanding of the methods to apply and their comparative detection power. Therefore, this paper presents an extensive and unique literature review, based on 97 papers from Web of Science and Scopus, resulting in a taxonomy following a recently proposed fraud analytics framework. We conclude that most research relies on expert-based rules and manual features, while deep learning methods have been gaining traction. This paper also presents a comprehensive framework to evaluate and compare the performance of prominent NA methods in a standardized setup. We apply it on two publicly available data sets, comparing manual feature engineering, random walk-based, and deep learning methods. We conclude that (1) network analytics increases the predictive power, but caution is needed when applying GNNs based on the class imbalance and network topology, and that (2) care should be taken with open-source data as this can give overly optimistic results. The open-source implementation facilitates researchers and practitioners to extend upon the results and experiment on proprietary data, promoting a standardized approach for the analysis and evaluation of network analytics for AML.
@@ 1 Introduction
Money laundering is the process of concealing illegally obtained funds by passing them through a complex sequence of transactions. Laundered money can be used in the legitimate financial system to fund and facilitate criminal activities, such as drug trafficking, corruption, and terrorism [77]. Around to of global GDP is laundered worldwide, amounting yearly to USD 2 trillion, as estimated by the United Nations Office on Drugs and Crime (UNODC) [130].
@@
Money laundering differs from other forms of financial fraud, especially credit card and payment fraud, as it is often done over a longer period and only becomes visible when looking at multiple transactions. It is more actively concealed, since sustained laundering is needed to process continuous streams of money. Credit card fraud, on the other hand, is much more about getting as much money out as fast as possible. Hence, the specific characteristics of money laundering merit dedicated research attention.
@@
Since, this illegally obtained money needs to enter the financial system at some point [77], legislative bodies have put stringent rules on financial institutions [15], requiring them to have safeguards and to report suspicious transactions/activities to the Financial Intelligence Unit (FIU). Despite record fines and mounting pressure by regulators, money laundering represents a growing problem [140]. Anti-money laundering practices are largely ineffective [35], with Europol estimating only 1.1% of criminal profits in the EU being confiscated [38].
Money laundering generally requires three main steps [130, 77]. The first is placement, where dirty money is integrated into the financial system. The second is layering, where money is transferred and mixed to obscure its illegal origin. The third is integration, where the laundered money is integrated into wider circulation by doing seemingly legitimate purchases.
Anti-money laundering practices are focused on placement and layering, as after integration via legitimate purchases, the money is in essence successfully laundered. Know-Your-Customer (KYC) protocols screen novel clients when they want to do business with the bank. Afterwards, transactions are continuously monitored to see if a client is involved in money laundering.
Anti-money laundering practices are often a combination of rule-based and risk-based systems [57, 123]. Rule-based systems consist of a collection of different IF-THEN rules. These are defined based on regulatory requirements and expert knowledge [110]. In practice, rule-based systems are still widely used [41, 22, 110]. The main drawback is that many rules are needed as money laundering is complex, which are hard to maintain. Risk-based systems apply more sophisticated methods. Past literature has shown that popular data analytics and machine learning methods have been applied for anti-money laundering [100, 57, 15, 76]. However, Le Khac and Kechadi [76] claim that most commercial tools still rely on more simple statistics to find outliers.
Placement, layering and integration involve multiple parties, and individual transactions often appear normal [15]. Therefore, network analytics is required to uncover the specific money laundering characteristics [15, 10]. The need for network analytics was already mentioned three decades ago by Senator et al. [116], where the authors used visualisations of relations to support experts investing suspicious cases. Although Ngai et al. [100] said that there was still a lack of research at that time, more attention has been given to money laundering and network analytics in the past years.
@@
The increasing attention has resulted in a growing body of literature covering network analytics for anti-money laundering. The swift expansion of the field has simultaneously resulted in a lack of a comprehensive overview. Most literature reviews on financial fraud barely discuss anti-money laundering using network analytics [100, 75, 42, 56]. Therefore, as a first step, this paper aims to supplement the literature by providing a systematic literature review of network analytics (NA) for anti-money laundering. To this end, we constructed queries to identify relevant papers from Web of Science (WOS) and Scopus, resulting in 97 papers.
The papers are analysed according to different dimensions. First, we look at the evolution of the number of published papers over time. Second, each paper is classified according to seven categories; publication data, the method type, the modelling method, the evaluation metrics used, the main objective, the kind of data, and network characteristics. We identify an increasing focus on crypto currencies. Our work compares the differences between those papers and the research covering classic financial transactions. Third, the top-cited work is reviewed via the framework introduced by Bockel-Rickermann et al. [14].
Our literature review and subsequent analysis reveals that multiple supervised, unsupervised and visualisation methods have been developed and applied. Nevertheless, due to the literature being fragmented and these methods mostly addressing only a part of the AML problem, there is limited insight into how different methods compare and which work best. Furthermore, most state-of-the-art (SOTA) network analytics methods have not yet found main-stream adoption for anti-money laundering. Therefore, this work provides a comprehensive experiment, evaluating the most prominent methods.
@@
We provide an extensive and critical systematic literature review on network analytics for anti-money laundering, to gain both comprehensive and deep insights, and provide direction for future research.
Based on our findings, we provide insights into the specific challenges of network methods based on the topology of the network and the data generating process.
@@
This paper is structured as follows. We present the methodology for the structured literature review in Section 2, with Section 3 presenting the results of the review. Section 4 describes the set-up of the empirical evaluation, with the results presented in Section 5. Finally, overall conclusions and directions for future work are presented in Section 6.
@@ 2.1 Methodology
We conducted a systematic literature search using the queries “graph analy*” AND “money launder*” and “network analy*” AND “money launder*” on WoS and Scopus, chosen for their high completeness and research credibility [97, 14]. The search is limited to English-language papers published before 2023, excluded recent publications to ensure adequate time for community evaluation.
The selection process included a title and abstract scan, followed by a full-text review. Papers were excluded if they focused on legal aspects or lacked substantial application of network analytics to anti-money laundering (AML). The final set of papers was classified according to seven criteria: publication data, method type, modelling approach, evaluation metrics, main objectives, data type, and network characteristics.

### v4

@@ Abstract
Money laundering presents a pervasive challenge, burdening society by financing illegal activities. The use of network information is increasingly being explored to effectively combat money laundering, given it involves connected parties. This led to a surge in research on network analytics for anti-money laundering (AML). The literature is, however, fragmented and a comprehensive overview of existing work is missing. This results in limited understanding of the methods to apply and their comparative detection power. This paper presents an extensive and unique literature review, based on 97 papers from Web of Science and Scopus, resulting in a taxonomy following a recently proposed fraud analytics framework. We conclude that most research relies on expert-based rules and manual features, while deep learning methods have been gaining traction. This paper also presents a comprehensive framework to evaluate and compare the performance of prominent methods in a standardized setup. We compare manual feature engineering, random walk-based, and deep learning methods on two publicly available data sets. We conclude that (1) network analytics increases the predictive power, but caution is needed when applying GNNs in the face of class imbalance and network topology, and that (2) care should be taken with synthetic data as this can give overly optimistic results. The open-source implementation facilitates researchers and practitioners to extend this work on proprietary data, promoting a standardized approach for the analysis and evaluation of network analytics for AML.
@@ 1 Introduction
Money laundering is the process of concealing illegally obtained funds by passing them through a complex sequence of transactions, so the money can be used to fund further criminal activities, e.g., drug trafficking, and terrorism [75]. The United Nations Office on Drugs and Crime [126] estimates that around to of global GDP is laundered worldwide, amounting yearly to USD 2 trillion.
@@
In contrast to credit card fraud, money laundering occurs over a longer duration and requires analysis of multiple transactions for detection. It is actively concealed, since sustained laundering is needed to process continuous money streams. Credit card fraud, on the other hand, aims to get as much money out as fast as possible. Hence, the specific characteristics of money laundering merit dedicated research attention.
@@
Since this money needs to enter the financial system at some point [75], legislators have put stringent rules on financial institutions [15] for reporting suspicious transactions. Despite record fines and mounting pressure by regulators, money laundering represents a growing problem [135].
Money laundering typically involves three stages [126, 75]: placement, where illicit funds enter the financial system; layering, where transactions obscure their origin; and integration, where the funds re-enter the economy through seemingly legitimate means, completing the laundering process.
Anti-money laundering practices require constant monitoring of clients and transactions. These practices are, however, largely ineffective [35], with Europol estimating only 1.1% of criminal profits in the EU being confiscated [38].
Effective money laundering involves multiple parties, with individual transactions appearing normal [15]. Therefore, network analytics is essential [15, 10]. This need was already mentioned three decades ago by Senator et al. [113], where the authors used visualisations to support expert investing.
@@
The increasing attention of the past years has resulted in a growing body of literature on network analytics for anti-money laundering, but simultaneously resulted in a lack of comprehensive overviews. Therefore, as a first step, this paper aims to supplement the literature by providing a comprehensive literature review of network analytics (NA) for anti-money laundering, covering 97 papers analysed according to multiple dimensions.
Due to the fragmented nature of the literature there is limited insight into which methods perform best. This fragmentation and lack of comparison is apparent in the fact that many of the introduced methods are not tested against any baselines. Furthermore, researchers are slow to adopt the latest network analytics methods and graph neural network in anti-money laundering. Therefore, the second goal of this work is to present a structured experimental set-up to evaluate the most prominent methods.
@@
We provide an extensive and critical literature review on network analytics for anti-money laundering, to gain comprehensive and deep insights, and provide directions for future research.
We benchmark a range of state-of-the-art network analytics methods for anti-money laundering, comparing the performance on two open-source data sets.
We provide insights into the specific challenges of network methods based on the topology of the network and the data generating process.
@@
This paper is structured as follows. We present the methodology for the literature review in Section 2, with Section 3 presenting the results and analysis. Section 4 describes the the empirical evaluation, with the results presented in Section 5. Overall conclusions and directions for future work are presented in Section 6.
@@ 2 Literature Review
The first part of this work covers the current literature. A summary is given according to multiple categories. Additional analyses are provided for the top-cited papers. The aim is to provide answers to the following research questions:
What are the most prominent learning methods in the literature and how have these evolved over time?
How are methods evaluated, using what data sets and evaluation metrics?
What is the setting, objectives and challenges discussed in the top-cited papers?
@@ 2.1 Methodology
We conducted the literature search using the queries “graph analy*” AND “money launder*” and “network analy*” AND “money launder*” on WoS and Scopus, chosen for their high completeness and research credibility [96, 14]. The search is limited to English-language papers published before 2023, excluded recent publications to ensure adequate time for community evaluation.
The selection process included a title and abstract scan, followed by a full-text review. Papers were excluded if they focused on legal aspects or lacked substantial application of network analytics to anti-money laundering (AML). The final set of papers is classified according to seven criteria as summarised in Table 1: publication data, method type, modelling approach, evaluation metrics, main objectives, data type, and network characteristics.